# Contact Wasserstein Geodesics for Non-Conservative Schrödinger Bridges

### v2

@@ C.1 Hamiltonian Structure of the Generalized Schrödinger Bridge
The optimality conditions resulting from the extremization of the cost functional $J$ in equation (15) follow from the Euler–Lagrange equations, generalized to the setting of classical field theory (blohmann2023lagrangian). In this framework, the arguments of the Lagrangian $\mathcal{L}$, in equation (16), are viewed as smooth fields defined over space and time. By setting to zero the variations of $\mathcal{L}$ with respect to these fields, we obtain the stationarity conditions for $J$. For a generic field $\psi^{t}(x)$, the corresponding Euler–Lagrange equation takes the form,
@@
The Hamiltonian system (19) can be reformulated in a linear and decoupled form via the coordinate transformation from the original variables $(\rho^{t},S^{t})$ to an alternative canonical set $(\hat{\rho}^{t},\hat{S}^{t})$. This transformation, known as Hopf–Cole coordinate transformation (leger2021hopf; chow2020wasserstein), is derived from the generating function $F$,
@@
and preserve the symplectic form, i.e., $dF=Sd\rho-\hat{S}d\hat{\rho}$. Following the standard generating-function rules (Chapter 18.3 10.1093/acprof:oso/9780198567264.001.0001),
@@
The Hopf–Cole transformation is well established in the literature, not only for simplifying the mathematical form of the equations, but also for enabling the design of efficient numerical integration schemes (leger2021hopf). In our context, it is particularly advantageous for obtaining a Hamiltonian with separable kinetic and potential energy components. The Hamiltonian system (19) then becomes,
@@
In this formulation, the Fisher information contributes to the potential energy and encodes the effect of stochastic diffusion. Minimizing the Fisher information term promotes smoothness in the density and steers the Hamiltonian flow toward the target distribution $\rho^{1}$, providing greater robustness due to the regularizing effect of diffusion. This mechanism has been studied in the literature and employed in control applications for its regularization properties (chen2025fictsc).
C.2 Contact Hamiltonian Structure of the Non-Conservative GSB
Here we derive the contact Hamiltonian formulation of the NCGSB problem, introduced in equation (4) and discussed in section 4. The structure of the derivation closely mirrors that of the GSB in Appendix C.1, with one key distinction: the Lagrangian $\mathcal{L}$ now depends explicitly on the accumulated action $z^{t}$. Specifically, the augmented cost functional and Lagrangian are given by,
@@ C.2 Contact Hamiltonian Structure of the Non-Conservative GSB
which induces the following coordinate change according to the contact generating rules in struckmeier2008covariant,
@@
C.3 Guided Schrödinger Bridge
@@
while preserving the underlying data manifold. Conditioning in this way modifies the probability flow $\rho^{t}$ (guo2024gradient) as,
$\rho^{t}(x\mid y)=\frac{1}{Z}\,\rho^{t}(x)\,e^{-\|y-f(\bar{x}^{t_{s}})\|^{2}},$ (42)
where $Z$ is a normalization constant, and $\bar{x}^{t_{s}}$ denotes a sample from the predicted prescribed distribution $\bar{\rho}^{t_{s}}$ conditioned on the current distribution $\rho^{t}$. This weight biases the generation toward samples that satisfy the desired property $y$. In a dynamical setting, we perform this conditioning by incorporating a control term $G^{t}$ into the Fokker–Planck dynamics,
@@
$G^{t}(x)=\nabla_{x^{t}}\log e^{-\|y-f(\bar{x}^{t_{s}})\|^{2}}=-\,\nabla_{x^{t}}\|y-f(\bar{x}^{t_{s}})\|^{2}.$ (45)
Specifically, in the NCGSB framework (4), this guidance is incorporated directly into the drift $v^{t}$ by adding the penalty $\|y-f(\bar{x}^{t_{s}})\|^{2}$ to the Lagrangian in the action dynamics constraint (4b),
$\begin{split}\min_{v^{t}}\quad&J(v^{t},\rho^{t})=\int_{0}^{1}\partial_{t}z^{t}\,dt,\\[3.00003pt] \text{s.t.}\quad&\partial_{t}z^{t}=\int_{\mathcal{M}}\left(\frac{1}{2}\|v^{t}(x)\|^{2}+U(x)+\|y-f(\bar{x}^{t_{s}})\|^{2}\right)\rho^{t}(x)\,dx-z^{t},\\[3.00003pt] &\partial_{t}\rho^{t}(x)+\nabla_{x}\cdot\big(\rho^{t}(x)\,v^{t}(x)\big)=\varepsilon\Delta_{x}\rho^{t}(x),\\[3.00003pt] &\rho^{0}=\rho_{a},\quad\rho^{1}=\rho_{b},\quad\rho^{t_{m}}=\rho_{m},\quad\forall\,m\in\{1,\dots,M\}.\end{split}$ (46)
Here, the inclusion of $\|y-f(\bar{x}^{t_{s}})\|^{2}$ in the Lagrangian produces the desired $\nabla_{x^{t}}\|y-f(\bar{x}^{t_{s}})\|^{2}$ correction in the drift, while preserving the Schrödinger bridge structure and constraints.
Appendix D ResNet Parameterization for Discrete Geodesics
D.1 Geometric Interpretation
@@ D.2 Training Algorithm
1:Dataset: samples from marginals $x_{a,\,\{i,j\}}\sim{\rho}_{a},\ x_{b,\,\{i,j\}}\sim{\rho}_{b},\ x_{n,\,\{i,j\}}\sim\{{\rho}_{n}\}_{n=1}^{N}$.
@@
16:   $\min_{\theta\setminus\theta^{0}}\ \ell(\{x^{t_{k}}\}_{k=0}^{K},x_{b,\,\{i,j\}},x_{n,\,\{i,j\}})$ $\triangleright$ Minimize geodesic loss (equation (9))
@@
D.3 Geodesic Parameterization: Time and Arc-Length Perspectives
@@
corresponds to a uniform arc-length discretization, where each segment has approximately the same length, i.e., $d^{2}_{\mathcal{W}_{2}}(\rho^{t_{k}}_{\theta},\rho^{t_{k-1}}_{\theta})\approx C$. The relationship between this arc-length discretization and the time discretization, which describes the progression along the geodesic with respect to the time variable $t$, is given by $ds=\sqrt{\mathcal{K}_{\text{phy}}}\,dt$, where $\mathcal{K}_{\text{phy}}$ denotes the kinetic energy of the system whose dynamics is interpreted as a geodesic. In the energy-conserving case, $\mathcal{K}_{\text{phy}}$ remains constant, and the time and arc-length discretizations differ only by a constant rescaling of the geodesic measure. Nevertheless, the discretization nodes $\{\rho_{\theta}^{t_{k}}\}_{k=0}^{K}$ remain unchanged, uniformly distributed in space and in time. In contrast, when modeling an energy-varying system whose associated trajectory projects onto a geodesic, the curve energy takes the form,
@@
where the scaling factor $\Phi^{t_{k}}=H(\rho^{t_{k}},S^{t_{k}},z^{t_{k}})-\mathcal{F}(\rho^{t_{k}})-{\mathcal{B}(\rho^{t_{k}})}$ accounts for the use of the Jacobi metric. This scaling factor also determines the length of each segment, which is inversely proportional to it, i.e., $d^{2}_{\mathcal{W}_{2}}(\rho^{t_{k}}_{\theta},\rho^{t_{k-1}}_{\theta})\approx\frac{C}{\Phi^{t_{k}}}$. Consequently, the arc-length discretization $d_{\mathcal{W}_{2}}(\rho^{t_{k}}_{\theta},\rho^{t_{k-1}}_{\theta})=\Delta s$ is non-uniform, while the time discretization $\Delta t=\sqrt{\frac{C}{\kappa}}$ remains uniform, as the nodes are assumed to be sampled at constant time intervals. The origin of this discrepancy lies in the kinetic energy term relating arc-length and time increments, $\Delta s=\sqrt{\mathcal{K}_{\text{phy}}}\Delta$, which is no longer constant in the energy-varying case.
This reasoning provides insight into the relationship between the energy term $\Phi^{t_{k}}$ used in the geodesic minimization and the physical kinetic energy of the modeled system. Specifically, since the geodesic minimization problem yields
@@
D.4 Time Complexity and Practical Considerations
At each iteration of the geodesic optimization of Alg. 1, we sample $N$ points $x^{s}\in\mathcal{M}$ of dimension $D$ from the reference distribution $\lambda$ and pass them through the ResNet. Assuming its architecture consists of $K$ blocks, each being an MLP of $L$ layers and hidden dimension $W$, the computational cost of this forward pass is $\mathcal{O}(N\,D\,K\,L\,W)$. The loss function $\ell$ in equation (9) requires $M+K$ evaluations of the Wasserstein distance between sample batches, and $K$ evaluations of the factor $\Phi^{t_{k}}$. For the Wasserstein distance, we employ the geomloss library (feydy2020geometric), which uses the Sinkhorn algorithm with time complexity $\mathcal{O}(N(D+T_{\text{sh}}))$, where $T_{\text{sh}}$ is the number of Sinkhorn iterations until convergence (feydy2020geometric). To evaluate $\Phi^{t_{k}}$, we apply its complete definition,

### v3

@@ D.1 Hamiltonian Structure of the Generalized Schrödinger Bridge
The optimality conditions resulting from the extremization of the cost functional $J$ in equation (15) follow from the Euler–Lagrange equations, generalized to the setting of classical field theory (Blohmann, 2024). In this framework, the arguments of the Lagrangian $\mathcal{L}$, in equation (16), are viewed as smooth fields defined over space and time. By setting to zero the variations of $\mathcal{L}$ with respect to these fields, we obtain the stationarity conditions for $J$. For a generic field $\psi^{t}(x)$, the corresponding Euler–Lagrange equation takes the form,
@@
The Hamiltonian system (19) can be reformulated in a linear and decoupled form via the coordinate transformation from the original variables $(\rho^{t},S^{t})$ to an alternative canonical set $(\hat{\rho}^{t},\hat{S}^{t})$. This transformation, known as Hopf–Cole coordinate transformation (Léger and Li, 2021; Chow et al., 2020), is derived from the generating function $F$,
@@
and preserve the symplectic form, i.e., $dF=Sd\rho-\hat{S}d\hat{\rho}$. Following the standard generating-function rules (Chapter 18.3 Johns (2005)),
@@
The Hopf–Cole transformation is well established in the literature, not only for simplifying the mathematical form of the equations, but also for enabling the design of efficient numerical integration schemes (Léger and Li, 2021). In our context, it is particularly advantageous for obtaining a Hamiltonian with separable kinetic and potential energy components. The Hamiltonian system (19) then becomes,
@@
In this formulation, the Fisher information contributes to the potential energy and encodes the effect of stochastic diffusion. Minimizing the Fisher information term promotes smoothness in the density and steers the Hamiltonian flow toward the target distribution $\rho^{1}$, providing greater robustness due to the regularizing effect of diffusion. This mechanism has been studied in the literature and employed in control applications for its regularization properties (Chen et al., 2025).
D.2 Contact Hamiltonian Structure of the Non-Conservative GSB
Here we derive the contact Hamiltonian formulation of the NCGSB problem, introduced in equation (4) and discussed in section 4. The structure of the derivation closely mirrors that of the GSB in Appendix D.1, with one key distinction: the Lagrangian $\mathcal{L}$ now depends explicitly on the accumulated action $z^{t}$. Specifically, the augmented cost functional and Lagrangian are given by,
@@ D.2 Contact Hamiltonian Structure of the Non-Conservative GSB
which induces the following coordinate change according to the contact generating rules in Struckmeier and Redelbach (2008),
@@
D.3 Guided Schrödinger Bridge
@@
while preserving the underlying data manifold. Conditioning in this way modifies the probability flow $\rho^{t}$ (Guo et al., 2024) as,
$\rho^{t}(x\mid y)=\frac{1}{Z}\,\rho^{t}(x)\,e^{-\|y-f({x}^{t_{s}})\|^{2}},$ (42)
where $Z$ is a normalization constant, and ${x}^{t_{s}}$ denotes a sample from the predicted prescribed distribution ${\rho}^{t_{s}}$(x—y). This weight biases the generation toward samples that satisfy the desired property $y$. In a dynamical setting, we perform this conditioning by incorporating a control term $G^{t}$ into the Fokker–Planck dynamics,
@@
$G^{t}(x)=\nabla_{x^{t}}\log e^{-\|y-f({x}^{t_{s}})\|^{2}}=-\,\nabla_{x^{t}}\|y-f({x}^{t_{s}})\|^{2}.$ (45)
Specifically, in the NCGSB framework (4), this guidance is incorporated directly into the drift $v^{t}$ by adding the penalty $\|y-f({x}^{t_{s}})\|^{2}$ to the Lagrangian in the action dynamics constraint (4b),
$\begin{split}\min_{v^{t}}\quad&J(v^{t},\rho^{t})=\int_{0}^{1}\partial_{t}z^{t}\,dt,\\[3.00003pt] \text{s.t.}\quad&\partial_{t}z^{t}=\int_{\mathcal{M}}\left(\frac{1}{2}\|v^{t}(x)\|^{2}+U(x)+\|y-f({x}^{t_{s}})\|^{2}\right)\rho^{t}(x)\,dx-z^{t},\\[3.00003pt] &\partial_{t}\rho^{t}(x)+\nabla_{x}\cdot\big(\rho^{t}(x)\,v^{t}(x)\big)=\varepsilon\Delta_{x}\rho^{t}(x),\\[3.00003pt] &\rho^{0}=\rho_{a},\quad\rho^{1}=\rho_{b},\quad\rho^{t_{m}}=\rho_{m},\quad\forall\,m\in\{1,\dots,M\}.\end{split}$ (46)
Here, the inclusion of $\|y-f({x}^{t_{s}})\|^{2}$ in the Lagrangian produces the desired $\nabla_{x^{t}}\|y-f({x}^{t_{s}})\|^{2}$ correction in the drift, while preserving the Schrödinger bridge structure and constraints.
Appendix E ResNet Parameterization for Discrete Geodesics
E.1 Geometric Interpretation
@@ E.2 Training Algorithm
1:Dataset: samples from marginals $x_{a,\,\{i,j\}}\sim{\rho}_{a},\ x_{b,\,\{i,j\}}\sim{\rho}_{b},\ x_{m,\,\{i,j\}}\sim\{{\rho}_{m}\}_{m=1}^{N}$.
@@
16:   $\min_{\theta\setminus\theta^{0}}\ \ell(\{x^{t_{k}}\}_{k=0}^{K},x_{b,\,\{i,j\}},x_{m,\,\{i,j\}})$ $\triangleright$ Minimize geodesic loss (equation (9))
@@
E.3 Stability Considerations
The proposed solver discretizes the geodesic flow on the Wasserstein manifold and can be interpreted as a single-shooting direct method for solving the optimal control problem of determining the optimal set of parameters $\{{\theta^{k}}^{*}\}_{k=0}^{K}$. Specifically, the ResNet-based parametrization $T_{\{\theta^{k}\}_{k=0}^{K}}$ produces a discrete sequence of intermediate batches $\{x^{t_{k}}_{\theta}\}_{k=0}^{K}$, converting the continuous optimal control problem into a nonlinear programming problem. This problem is then solved using standard static optimization methods, with gradients computed through backpropagation. The loss function in equation 9,
$\ell=\underbrace{d_{\mathcal{W}_{2}}^{2}({\rho}_{\theta}^{t_{K}},{\rho}_{b})}_{\text{Terminal marginal}}+\underbrace{\sum_{m=1}^{M}d_{\mathcal{W}_{2}}^{2}({\rho}_{\theta}^{t_{k_{m}}},{\rho}_{m})}_{\text{Intermediate marginals}}+\underbrace{\sum_{k=1}^{K}\Phi^{t_{k}}\,d_{\mathcal{W}_{2}}^{2}({\rho}_{\theta}^{t_{k}},{\rho}_{\theta}^{t_{k-1}})}_{\text{Energy minimization}},$
involving the computation of the Wasserstein distance, which can be explicitly rewritten as,
$\ell=\underbrace{\sum_{i,l}^{N}\pi_{il}\|x_{i}^{t_{K}}-y_{l}^{t_{K}}\|^{2}}_{\text{Terminal marginal}}+\underbrace{\sum_{m=1}^{M}\sum_{i,l}^{N}\pi_{il}\|x_{i}^{t_{k_{m}}}-y_{l}^{t_{m}}\|^{2}}_{\text{Intermediate marginals}}+\underbrace{\sum_{k=0}^{K-1}\sum_{i,l}^{N}\Phi^{k}\pi_{il}\|x_{i}^{t_{k+1}}-x_{i}^{t_{k}}\|^{2}}_{\text{Energy minimization}},$ (47)
where $\pi_{il}$ denotes the optimal transport plan between samples $i$ and $l$ the data batches, optimized in order to minimize the transport cost and preserve the constraint, $\sum_{il}^{N}\pi_{il}=1$, ensuring full mass conservation between batches. Importantly, the optimization of the transport plan $\pi_{il}$ is a subproblem carried out for a fixed set of parameters $\{\theta^{k}\}_{k=0}^{K}$. This subproblem is a linear program whose solution depends Lipschitz-continuously on $\{\theta^{k}\}_{k=0}^{K}$, i.e.,
$\left\|\pi\left({\{\theta^{k^{\prime}}\}_{k=0}^{K}}\right)-\pi\left({\{\theta^{k}\}_{k=0}^{K}}\right)\right\|\leq L_{\pi}\left\|{\{\theta^{k^{\prime}}\}_{k=0}^{K}}-{\{\theta^{k}\}_{k=0}^{K}}\right\|,$ (48)
where $L_{\pi}$ is a scalar constant. The two differences quantify, respectively, the change in the transport plan induced by a perturbation in the parameter set, and the perturbation in the parameter set itself.
The computation of the gradient of the loss in equation (47), which is responsible for updating the set of parameters according to,
$\{\theta^{k}_{(j+1)}\}_{k=0}^{K}=\{\theta^{k}_{(j)}\}_{k=0}^{K}-\alpha\,\nabla_{\theta}\ell,$ (49)
leads to the appearance of three types of terms:
  • $\nabla_{\theta}\pi_{il}$, which is Lipschitz in $\{\theta^{k}\}_{k=0}^{K}$ because the transport plan $\pi$ is the solution of a convex optimization problem whose dependence on the parameters is Lipschitz-continuous;
  • $\nabla_{\theta}\|x_{i}^{t_{k}}-y_{l}^{t_{k}}\|^{2}$, where $x_{i}^{t_{k}}$ depends nonlinearly on $\{\theta^{k}\}_{k=0}^{K}$ due to the nonlinear ResNet parametrization $T_{\{\theta^{k}\}_{k=0}^{K}}$ used for expressivity. This implies that multiple different parameter sets may minimize this term. However, around a local optimal point the problem of minimizing the Euclidean distance is locally convex;
  • $\nabla_{\theta}\Phi^{t_{k}}$, which is a nonlinear function of $\{x^{t_{k}}_{\theta}\}_{k=0}^{K}$. If $\Phi^{t_{k}}$ is at least piece-wise smooth, it is possible to identify a local minimum for $\{x^{t_{k}}_{\theta}\}_{k=0}^{K}$, and therefore a corresponding local minimum for $\{\theta^{k}\}_{k=0}^{K}$, depending on the initialization of the parameters. Even in the most complex scenario considered in this paper, the energy-varying SB case, the required smoothness assumptions are satisfied. Specifically, for $\Phi^{t_{k}}=H(\rho^{t_{k}},S^{t_{k}},z^{t_{k}})+U(\rho^{t_{k}})+I(\rho^{t_{k}})-{\mathcal{B}(\rho^{t_{k}})}$, $H$ is a smooth function depending on a hyperparameter, $U$ is implemented either via at least piecewise smooth learning models (Apps. G.4, G.5, G.7) or via a LAND metric (Arvanitidis et al., 2018) with a smooth kernel (Apps. G.2, G.3), and both $I$ and $\mathcal{B}$ are smooth functions.
Therefore, each update $\{\theta^{k}_{(j)}\}_{k=0}^{K}\;\longrightarrow\;\{\theta^{k}_{(j+1)}\}_{k=0}^{K}$ does not destabilize the training and converges toward a local minimum, which geometrically corresponds to one of the possible coordinate charts that can describe the tangent bundle $\mathcal{T}\mathcal{P}^{+}(\mathcal{M})$ (see Appendix E.1 for the geometrical interpretation). The stability of the gradient for long time horizons is moreover guaranteed by the ResNet architecture itself, where the skip connections prevent detrimental effects such as exploding or vanishing gradients (Zaeemzadeh et al., 2020).
E.4 Geodesic Parameterization: Time and Arc-Length Perspectives
@@
corresponds to a uniform arc-length discretization $s$, where each segment has approximately the same length, i.e., $d^{2}_{\mathcal{W}_{2}}(\rho^{t_{k}}_{\theta},\rho^{t_{k-1}}_{\theta})\approx C$. The relationship between the arc-length discretization $s$ and the time discretization $t$, which describes the progression along the geodesic with respect to the time variable, is given by $ds=\sqrt{\mathcal{K}_{\text{phy}}}\,dt$, where $\mathcal{K}_{\text{phy}}$ denotes the kinetic energy of the system whose dynamics is interpreted as a geodesic. In the energy-conserving case, $\mathcal{K}_{\text{phy}}$ remains constant, and the time and arc-length discretizations differ only by a constant rescaling of the geodesic length. Nevertheless, the discretization nodes $\{\rho_{\theta}^{t_{k}}\}_{k=0}^{K}$ remain unchanged, uniformly distributed in space and in time. In contrast, when modeling an energy-varying system whose associated trajectory projects onto a geodesic, the curve energy takes the form,
@@
where the scaling factor $\Phi^{t_{k}}=H(\rho^{t_{k}},S^{t_{k}},z^{t_{k}})-\mathcal{F}(\rho^{t_{k}})-{\mathcal{B}(\rho^{t_{k}})}$ accounts for the use of the Jacobi metric. This scaling factor also determines the length of each segment, which is inversely proportional to it, i.e., $d^{2}_{\mathcal{W}_{2}}(\rho^{t_{k}}_{\theta},\rho^{t_{k-1}}_{\theta})\approx\frac{C}{\Phi^{t_{k}}}$. Consequently, the arc-length discretization $d_{\mathcal{W}_{2}}(\rho^{t_{k}}_{\theta},\rho^{t_{k-1}}_{\theta})=\Delta s$ is non-uniform, while the time discretization $\Delta t=\sqrt{\frac{C}{\kappa}}$ remains uniform, as the nodes are assumed to be sampled at constant time intervals. The origin of this discrepancy lies in the kinetic energy term relating the arc-length and time increments, $\Delta s=\sqrt{\mathcal{K}_{\text{phy}}}\Delta$, which is no longer constant in the energy-varying case.
This reasoning provides insight into the relationship between the energy term $\Phi^{t_{k}}$ used in the geodesic minimization and the physical kinetic energy of the modeled system. Specifically, since the geodesic minimization problem yields,
@@
E.5 Time Complexity and Practical Considerations
At each iteration of the geodesic optimization of Alg. 1, we sample $N$ points $x^{s}\in\mathcal{M}$ of dimension $D$ from the reference distribution $\lambda$ and pass them through the ResNet. Assuming its architecture consists of $K$ blocks, each being an MLP of $L$ layers and hidden dimension $W$, the computational cost of this forward pass is $\mathcal{O}(N\,D\,K\,L\,W)$. The loss function $\ell$ in equation (9) requires $M+K$ evaluations of the Wasserstein distance between sample batches, and $K$ evaluations of the factor $\Phi^{t_{k}}$. For the Wasserstein distance, we employ the geomloss library (Feydy, 2020), which uses the Sinkhorn algorithm with time complexity $\mathcal{O}(N(D+T_{\text{sh}}))$, where $T_{\text{sh}}$ is the number of Sinkhorn iterations until convergence (Feydy, 2020). To evaluate $\Phi^{t_{k}}$, we apply its complete definition,